OpenStudy (anonymous):

3|x − 8| + 15 = 18 A. x = 1, x = 17 B. x = −7, x = 9 C. x = 7 x = 9 D. No solution kinda confused on this question, help please?

OpenStudy (irishboy123):

3|x − 8| + 15 = 18 3|x − 8| = 3 ie proceed with normal algebra until you arrive at |x − 8| = k then you know that the solutions are: x − 8 = k, AND x − 8 = -k and on you go to solve for x